# Barking Up The Syntactic Tree: Enhancing VLM Training with Syntactic Losses

### v1

@@ 4.2 Complex Referring Segmentation
As shown in Table 3, we outperform both ALBEF and SelfEQ [14] by 6.2% and 5.3%, respectively, on the complex scene referring segmentation. Additionally, Figure 3 shows qualitative comparisons among ALBEF, ALBEF + SelfEQ, and ALBEF + HIST. We observe that ALBEF + HIST significantly outperforms ALBEF and ALBEF + SelfEQ in indentifying complex queries involving multiple objects within intricate scenes. For instance, it accurately identifies the “Blurry pie” in a blurry image, the occluded “Not full zebra” and relative object positions such as “The further meter,” the “Nearest meter,” and the “Lady in the back.” This enhanced ability to distinguish multiple objects in challenging contexts underscores HIST’s robustness and effectiveness in visual grounding tasks. (More qualitative result provided in the supplementary.)
@@ 4.3 Downstream Tasks
While our losses were designed to improve localization, we also test the impact of they have on other downstream VLM tasks, including image-text retrieval and VQA.
@@
ALBEF [24] 77.6 60.7 95.9 85.6
(Ours) ALBEF + HIST 78.7 61.4 96.1 86.0
@@
ALBEF [24] 75.8 76.0
(Ours) ALBEF + HIST 76.0 76.2
@@ 4.4 Ablation Study
HIST Losses. We ablate different loss components of HIST + ALBEF and report the results in Tab. 6. Each component contributes positively to the final performance, which highlights the significance of our design choices.
Data Objective Flickr30k
- $\mathcal{L}_{\text{vl}}$ 79.4
$T$ $\mathcal{L}_{\text{phrase}}$ 81.6
$T_{HIST}$ $\mathcal{L}_{\text{phrase}}$ + $\mathcal{L}_{\text{add}}$ 85.0
$T_{HIST}$ $\mathcal{L}_{\text{phrase}}$ + $\mathcal{L}_{\text{subj}}$ + $\mathcal{L}_{\text{add}}$ 85.3
Table 6: Ablation on HIST Losses. Models are trained with COCO. T represents training with only noun phrases of COCO. $T_{HIST}$ represents training with our HIST structure.
BLIP Layer and Head effect. As shown in Table 7, our experiments demonstrate that when training BLIP with HIST, using the head average attention map from the 9th cross-attention layer to calculate the Addition Loss enhances grounding performance in both layer 9 head 10, and layer 8 head 10 of BLIP. Consequently, we report BLIP’s visual grounding performance by training with the Addition loss using the head average Grad-CAM map from the 9th cross-attention layer and evaluate using the mean of two attention maps: (1) layer 9 head 10 and (2) layer 8 head 10.
Eval Train L8 L9 L8H10 L9H10 Mean Mul
L8H10 51.1 56.9 42.9 56.5 48.9 56.5
L9H10 48.3 51.8 53.9 32.4 17.1 32.4
Mean 51.2 57.6 53.9 56.7 18.7 56.8
Mul 52.4 56.2 53.7 49.6 21.4 49.6
Table 7: Ablation of HIST + BLIP on layer and head. We train HIST+BLIP on COCO and test on Referit with image size of 384. Mean represent the average of attention maps from L8H10 and L9H10 of BLIP. Mul represent the element-wise multiplication of attention maps from L8H10 and L9H10 of BLIP.
@@ 5 Conclusion
We present a novel HIerarchically STructured (HIST) learning framework that enhances vision-language alignment in Vision-Language Models (VLMs) by embedding hierarchical syntactic structures into training. HIST organizes captions into structured noun phrases across multiple syntactic levels, improving alignment at both phrase and object levels without additional supervision. Our framework introduces two key losses: Subject Loss, ensuring alignment with object-centric phrases and subjects, and Addition Loss, promoting balanced attention across multiple objects in complex scenes. HIST significantly enhances localization in challenging multi-object scenes and improves performance in tasks like image-text retrieval and visual question answering, underscoring the value of hierarchical structuring for spatial and semantic alignment in VLMs.

### v2

@@ 4.2 Complex Referring Segmentation
As shown in Table 4, we outperform both ALBEF and SelfEQ [16] by 6.2% and 5.3%, respectively, on the complex scene referring segmentation. Additionally, Figure 3 shows qualitative comparisons among ALBEF, ALBEF + SelfEQ, and ALBEF + HIST. We observe that ALBEF + HIST significantly outperforms ALBEF and ALBEF + SelfEQ in indentifying complex queries involving multiple objects within intricate scenes. For instance, it accurately identifies the “Blurry pie” in a blurry image, the relative object positions such as the “Lady in the back”, “Glass in between mugs” and “Food nearest to us”. This enhanced ability to distinguish multiple objects in challenging contexts underscores HIST’s robustness and effectiveness in visual grounding tasks.
(More qualitative result provided in the supplementary.)
Method Training CIoU
[…]
        + HIST(Ours) COCO 35.6
Table 4: Complex Referring Segmentation on RefCOCO+ [58].
@@ 4.3 Downstream Tasks
While our losses are designed to enhance localization, we also evaluate their impact on other downstream VLM tasks, including image-text retrieval and VQA.
@@
[…]
        + HIST(Ours) 78.7 61.4 96.1 86.0
@@
[…]
        + HIST(Ours) 76.0 76.2
@@ 4.4 Ablation Study
HIST Losses. We ablate different loss components of HIST applied to ALBEF and report the results in Table 7. Each component contributes positively to the final performance, which highlights the significance of our design choices. Importantly, Composition accounts for most significant improvement which goes to our core technical contribution.
Data Objective Flickr30k Referit
- $\mathcal{L}_{\text{vl}}$ 79.4 59.7
$T$ $\mathcal{L}_{\text{phrase}}$ 73.2 63.2
$T_{HIST}$ $\mathcal{L}_{\text{phrase}}$ + $\mathcal{L}_{\text{subj}}$ 75.0 64.0
$T_{HIST}$ $\mathcal{L}_{\text{phrase}}$ + $\mathcal{L}_{\text{comp}}$ 83.2 69.4
$T_{HIST}$ $\mathcal{L}_{\text{phrase}}$ + $\mathcal{L}_{\text{comp}}$ + $\mathcal{L}_{\text{subj}}$ 83.6 69.5
Table 7: Ablation on HIST Losses. Models are trained with VG. T represents training with only phrases of VG. $T_{HIST}$ represents training with our HIST structure.
BLIP Layer and Head effect. From our ablation experiments (detail results in Sec 8 of supplementary), when training BLIP with HIST, using the head average attention map from the 9th cross-attention layer to calculate the Composition Loss enhances grounding performance in both layer 9 head 10, and layer 8 head 10 of BLIP. Consequently, we report BLIP’s visual grounding performance by training with the Composition loss using the head average Grad-CAM map from the 9th cross-attention layer and evaluate using the mean of two attention maps: (1) layer 9 head 10 and (2) layer 8 head 10.
5 HIST on TinyLLaVA
To show the generalization of our method, we apply HIST on TinyLLaVA 0.5B, a LLaVA-like variant, to demonstrate the benefit of hierarchical structured data on spatial image-text alignment. TinyLLaVA 0.5B has the same structure as LLaVA [33] but with SigLIP [61] as visual encoder and small-scale Qwen2-0.5B [38] as LLM.
Reformulate data to QA format. To adapt HIST on TinyLLaVA, we first reformulate the data into QA format that fits the language modeling loss. Specifically, we replace each phrase in the original caption with a special token [blank] and prompt the model to “fill in the [blank]” by outputting the original phrase. As an example, for a caption “a white cat and a gray dog playing on the ground”, the input to the model would be “[blank] and a gray dog playing on the ground. Please fill in the [blank].” and the answer would be “a white cat”. For the composite phrase, the input to the model would be “[blank] and [blank] playing on the ground. Please fill in the [blank].” and the answer would be “a white cat and a gray dog”.
Attention. To extract the attention of the pair of phrases $\mathbf{C}_{i}$ and $\mathbf{C}_{j}$ and the composite phrase $\mathbf{C}_{ij}$ for Composition loss, we make use of the self-attention in Qwen2. Qwen2 takes as input a sequence of length $N$ composed of (i) default text prompt tokens, (ii) $P\times P$ image tokens, (iii) input query tokens and (iv) $K$ output answer tokens, and output the corresponding next-token predictions. From specific layers, we take the $K\times(P\times P)$ answer-to-image sub-attention map $\mathbf{A}$ from the $N\times N$ self-attention map, take the mean over all $K$ answer tokens, and reshape the vector to obtain the $P\times P$ phrase specific map. With the corresponding maps of $\mathbf{A}_{\mathbf{C}_{i}}$, $\mathbf{A}_{\mathbf{C}_{j}}$, and $\mathbf{A}_{\mathbf{C}_{ij}}$, we calculate the Composition loss as in Eq. 5. The Composition loss is combined with the language modeling loss to finetune the TinyLLaVA model.
Visual Grounding Result. As shown in Table 8, compared to Tinyllava 0.5B finetuned on COCO with the original caption, Tinyllava + HIST outperform by 10.6% on Flickr30K and 5.0% on Referit. This demonstrates the general applicability of HIST on both contrastive and generative VLMs.
Figure 4: Qualitative Result for TinyLLaVA and TinyLLaVA + HIST on Visual Grounding. Images are from Flickr30K.
Method Training Flickr30K Referit
TinyLLaVA [65] COCO 36.4 26.0
       + HIST(Ours) COCO 47.0 31.0
Table 8: Visual Grounding results of HIST on TinyLLaVA with Pointing Game Accuracy as metric.
@@ 6 Conclusion
We present a novel HIerarchically STructured (HIST) learning framework that enhances vision-language alignment in Vision-Language Models (VLMs) by embedding hierarchical syntactic structures into training. HIST organizes captions into structured phrases across multiple syntactic levels, improving alignment at both phrase and object levels without additional annotation. Our framework introduces two key losses: Subject Loss, ensuring alignment with object-centric phrases and subjects, and Composition Loss, promoting balanced attention across multiple objects in complex scenes. HIST significantly enhances localization in challenging multi-object scenes and improves performance in tasks like image-text retrieval and visual question answering, underscoring the value of hierarchical structuring for spatial and semantic alignment in VLMs.